OpenStudy (anonymous):

What is the distance between the points (21, 16) and (9, 11)?

OpenStudy (anonymous):

\[\sqrt((21-9)^2 + (16-11)^2) = \sqrt(12^2 + 5^2) = \sqrt(169) = 13\]